OpenStudy (anonymous):

twelve people arrive at restaurant. there is one table for six, one table for 4 and one table for 2. In how many ways can they be assigned to a table

OpenStudy (anonymous):

good morning imran just did this problem, but i forget about the circle

OpenStudy (anonymous):

this isn't a circle problem though

OpenStudy (anonymous):

Since total Head count is 12. So for 1st table for six the probability is 6*12=72 Now only six is left and so for 2nd table for four the probability is 4*6=24 Now only two is left and so for 3rd table for 2 the probability is 2 so total no. of ways is 72+24+2=94 Is it correct?

OpenStudy (anonymous):

no that is not correct

OpenStudy (anonymous):

no the answer is 13 860

OpenStudy (anonymous):

what i did was : 12 C 6 + 12 C 4 + 12 C 2

OpenStudy (anonymous):

but its wrong lol

OpenStudy (anonymous):

oh wait i should learn to read. it doesn't say how many arrrangements are there, just how many possible assignments

OpenStudy (anonymous):

12 people 12 chairs

OpenStudy (anonymous):

the circle arrangement doesn't apply to this question imran

OpenStudy (anonymous):

should be \[\dbinom{12}{6}\times \dbinom{6}{4}\]

OpenStudy (anonymous):

cause it doesn't state that the tables are circuar

OpenStudy (anonymous):

no i said that and i was wrong

OpenStudy (anonymous):

can you explain it to me sattellite

OpenStudy (anonymous):

first off is the answer 13860?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

It does not say the tables are separated or put together just says a table that holds 6, 4 and 2 which would get your 12 chairs and you have 12 people now how many combinations can you create

OpenStudy (anonymous):

because i don't want to start explaining something that is wrong!

OpenStudy (anonymous):

lol no ur correct

OpenStudy (anonymous):

ok so we first have to be clear about what the problem is asking. it wants the number of ways you can partition 12 people in to one group of 6, one group of 4 and one group of 2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Ok so the tables are separated

OpenStudy (anonymous):

pretty sure they are

OpenStudy (anonymous):

out of 12 people you are going to choose 6 to be at table A. how many ways are there to do this? the answer is \[\dbinom{12}{6}\] which really just restates the question. we compute \[\dbinom{12}{6}=\frac{12\times11\times 10\times 9\times 8\times7}{6\times 5\times 4\times 3\times2}=924\]

OpenStudy (anonymous):

that is the number of ways to choose 6 from a set of 12

OpenStudy (anonymous):

now you have 6 people left and you want the number of ways to assign them to table B that is \[\dbinom{6}{4}=\frac{6\times 5}{2}=15\]

OpenStudy (anonymous):

notice you could have said "how many ways to assign them to table C?" and got \[\dbinom{6}{2}\] which is also 15

OpenStudy (anonymous):

now you are done because once you have assigned 6 to table A and 4 to table B you have no choices for table C

OpenStudy (anonymous):

so 924 ways to make up table A, 15 ways to make up table B, then only one way for table C and by counting principle the number of ways to do this all together is \[924\times15 =13860\]

OpenStudy (anonymous):

thanks so much for the explanation:)

OpenStudy (anonymous):

yw hope it is relatively clear